### v1

# Constraint-Aware Reinforcement Learning via Adaptive Action Scaling

Abstract
Safe reinforcement learning (RL) seeks to mitigate unsafe behaviors that arise from exploration during training by reducing constraint violations while maintaining task performance. Existing approaches typically rely on a single policy to jointly optimize reward and safety, which can cause instability due to conflicting objectives, or they use external safety filters that override actions and require prior system knowledge. In this paper, we propose a modular cost-aware regulator that scales the agent’s actions based on predicted constraint violations, preserving exploration through smooth action modulation rather than overriding the policy. The regulator is trained to minimize constraint violations while avoiding degenerate suppression of actions. Our approach integrates seamlessly with off-policy RL methods such as SAC and TD3, and achieves state-of-the-art return-to-cost ratios on Safety Gym locomotion tasks with sparse costs, reducing constraint violations by up to 126 times while increasing returns by over an order of magnitude compared to prior methods.
I Introduction
Reinforcement Learning (RL) has demonstrated remarkable success across a range of domains, including Atari games [1], robotics [2, 3], and long-horizon strategy games [4, 5]. This success is significantly facilitated by exploratory behavior, which allows agents to discover effective behaviors. However, such exploratory behaviors often lead to the violation of constraints imposed on the controlled system. While constraint violations are tolerable in simulated environments and games where resets are free, they pose serious risks in real-world applications [6]. Violating safety constraints can lead to irreversible damage or system failure. To address this issue, Safe Reinforcement Learning (Safe RL) [7] has emerged as a critical area of research that aims to minimize constraint violations during both training and deployment.
Safe RL methods can be broadly categorized into two groups: safe exploration and constrained RL. Safe exploration techniques aim to prevent the agent from taking actions that violate safety constraints. These methods typically rely on prior knowledge of the system dynamics and feasible safe states to construct control barrier functions [8, 9, 10], or model predictive shields [11, 12, 13]. Although effective, their applicability is limited by the need for detailed prior information about the system dynamics, an assumption that often does not hold in early learning stages or tasks where system dynamics are unknown.
Figure 1: Overview of cost-aware action scaling. The RL agent proposes an action that would result in the center of mass (COM) exceeding the velocity threshold (left). The regulator (blue) intervenes by scaling the action, keeping the velocity of the COM within the safe zone while allowing progress on the task. The yellow circles highlight the velocity threshold for the COM, illustrating how the regulator enforces a safety constraint while preserving task performance.
Constrained RL instead allows the agent to learn both reward and cost signals online, without requiring knowledge of the system dynamics. The agent is trained to maximize cumulative rewards while minimizing constraint violations. Common approaches include Lagrangian-based methods that solve the dual formulation of the constrained optimization problem [14, 15, 16, 17], and budget-based methods maintain a running estimate of remaining safety allowance, adjusting the agent’s behavior to respect cost limits over time [18, 19]. However, a core limitation of these methods is the difficulty of balancing reward and cost within a single policy. Conflicting gradients can cause the agent to behave either too conservatively or unsafely, leading to instability, constraint violations, or poor performance [17, 20].
In contrast to prior work, we propose a modular alternative: instead of overriding actions or jointly optimizing conflicting objectives, we scale actions based on the expected cost of future constraint violations while preserving the policy’s task-directed behavior (see Fig. 1).The architecture consists of a reward-maximizing task agent and a regulator network guided by twin cost critics to conservatively estimate constraint violations. The regulator applies element-wise scaling to attenuate risky actions, enforcing safety without requiring prior knowledge of dynamics or compromising exploration. Although our approach resembles safe exploration in formulation, we do not require prior knowledge of the system dynamics, and we do not override the task agent’s actions, thereby preserving both exploration and safety without external overrides.
We evaluate our approach on several dynamical systems from Safety Gymnasium [21] and the Safety-Critical environments from [22]. Our method achieves the highest Return-to-Cost (RC) ratio [23], reducing constraint violations by up to 126 times over recent safe RL baselines [17, 19, 24, 25, 26]. In summary, our contributions are:
  • We propose a modular safe RL framework that decouples reward maximization and safety enforcement via a cost-aware regulator that scales actions based on predicted violations.
  • Our model-free approach integrates seamlessly into standard off-policy RL pipelines such as SAC [27] and TD3 [28], improving safety without compromising exploration.
  • We achieve state-of-the-art performance on safety benchmarks, with up to 126 times fewer constraint violations and the highest RC ratios across tasks.
II Related Work
Safe RL aims to enable agents to maximize task rewards while minimizing constraint violations. Existing approaches broadly fall into two categories: safe exploration methods, which intervene externally to prevent unsafe actions during training, and constrained RL methods, which embed cost objectives directly into policy optimization.
Safe Exploration Methods. Safe exploration techniques intervene during training to prevent agents from entering unsafe states. Early methods, such as [29], employed uncertainty modeling through Gaussian Processes to restrict exploration and ensure safety during optimization. Later approaches introduced safety layers [30, 31] and predictive safety filters [11, 12, 13, 22] that anticipate and block risky actions based on pre-trained layers or model predictive control (MPC). Control Barrier Function (CBF)-based strategies [8, 9, 10] encode safety constraints directly through differentiable control barriers to guarantee that the agent’s actions remain within certified safe sets throughout training. [32] extend shielding methods to continuous domains by leveraging approximate dynamics models, enabling probabilistic safety guarantees during exploration. [33] leveraged model-based RL and offline collected data to develop reachability-based safety layers to ensure safe actions for navigation scenarios. [23, 34] assumes access to an offline dataset for pretraining a cost critic along with a recovery policy, which is then fixed during online learning, limiting its applicability in settings where collecting sufficient offline data is challenging or costly.
While our method shares some conceptual similarity with these safe exploration approaches—modifying actions to maintain safety—it differs fundamentally by relying on online-learned cost predictions rather than external models or handcrafted safe sets, and by smoothly scaling actions instead of hard blocking or overwriting them, preserving the agent’s exploratory behavior.
Constrained RL Approaches. Constrained RL methods integrate cost minimization into policy learning itself. Lagrangian-based algorithms [14, 16, 17] optimize dual formulation balancing rewards and costs, while budgeted RL approaches [18, 19] include the remaining cost budget in the state representation, allowing the agent to adapt its behavior based on how much cost it can still afford. Risk-sensitive formulations, such as CVaR-CPO [35], enforce safety by constraining the conditional value-at-risk of cumulative costs, ensuring attention to costly violations. Reachability-based methods like RESPO [25] estimate the probability of reaching safe regions and optimize policies to persistently satisfy constraints or recover when outside the feasible set. Constraint-Conditioned Policy Optimization [36] enables zero-shot generalization to unseen cost thresholds by conditioning the policy and value functions on constraint levels using a variational inference objective. Bi-level optimization frameworks such as SRCPO [26] address the nonlinearity of risk measures by optimizing over dual variables, achieving strong constraint satisfaction in continuous control tasks. Safety Editor [24] instead trains two separate Soft Actor-Critic (SAC) [27] agents: a utility maximizer and a safety editor that modifies unsafe actions, allowing it to fully overwrite the original action when necessary.
Compared to these methods, our approach offers a lightweight, modular alternative: instead of embedding constraints into the policy loss, relying on delicate dual updates, or training a second actor to overwrite unsafe actions, we regulate actions externally using learned cost estimators. This continuous scaling preserves exploration while enabling seamless integration into standard off-policy RL pipelines.
Figure 2: Overview of our modular safe RL architecture. The regulator (blue) scales actions produced by the unconstrained RL agent (yellow) based on predicted cost (purple), producing safety-aware actions (green) that are executed in the environment.
III Preliminaries
Markov Decision Processes. We consider a Markov Decision Process (MDP) defined by the tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,r,\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $P(s^{\prime}|s,a)$ the transition probability, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ the reward function, and $\gamma\in(0,1)$ the discount factor. We assume continuous state and action spaces with $\mathcal{S}\subseteq\mathbb{R}^{n}$ and $\mathcal{A}\subseteq\mathbb{R}^{d}$.
Constrained Reinforcement Learning. Constrained RL refers to the problem of maximizing task rewards while satisfying explicit constraints on agent behavior. The standard formalism is through Constrained Markov Decision Processes (CMDPs) [37], where constraints are expressed using cost functions separate from the reward.
Constrained Markov Decision Processes. A CMDP augments the MDP with a cost function $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}_{\geq 0}$ that quantifies safety violations. The objective is to maximize return while keeping the expected cumulative cost below a budget $\chi$:
$\max_{\pi}\;$ $\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot|s)}[Q^{\pi}(s,a)]$
s.t. $\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot|s)}[Q_{c}^{\pi}(s,a)]\leq\chi,$ (CMDP)
where the cost value functions are
$V_{c}^{\pi}(s)$ $=\mathbb{E}_{\pi}\Big[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\,\Big|\,s_{0}=s\Big],$ (1)
$Q_{c}^{\pi}(s,a)$ $=\mathbb{E}_{\pi}\Big[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\,\Big|\,s_{0}=s,\,a_{0}=a\Big].$ (2)
Cost Budget. The budget $\chi$ specifies the maximum allowable expected cumulative cost and is typically treated as a human-selected threshold that reflects task-specific safety requirements [17]. In this work, we assume a stricter setting by eliminating the cost budget, i.e., setting $\chi=0$, similar to [25]. This corresponds to a hard-safety regime that aims to achieve minimal constraint violations during learning.
Problem Setting. With this stricter formulation, the problem considered in this work is to learn a policy that maximizes task rewards while minimizing constraint violations under the hard-safety regime $\chi=0$. Formally, our objective reduces to:
$\pi^{*}$ $=\arg\max_{\pi}\;\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot|s)}[Q^{\pi}(s,a)]$ (3)
s.t. $\;\mathbb{E}_{s\sim d_{0},\,a\sim\pi(\cdot|s)}[Q_{c}^{\pi}(s,a)]=0.$ (4)
This assumption eliminates any positive cost budget and focuses on policies that aim to achieve minimal safety violations during training and execution.
IV Methodology
We propose a modular safe reinforcement learning framework that regulates the actions of a task policy to reduce expected constraint violations without overriding agent decisions. The key idea is to scale actions based on their predicted cost, preserving exploration while inducing smoother and safer transitions in the environment.
IV-A Split Architecture for Reward and Cost Optimization
[…]
IV-B Action Modulation via Regulator Scaling
In continuous control environments without stochasticity, the system evolves under deterministic transition dynamics of the form $s_{t+1}=f(s_{t},a_{t})$, where $s_{t}\in\mathcal{S}$ is the current state and $a_{t}\in\mathcal{A}\subset\mathbb{R}^{d}$ is a $d$-dimensional real-valued action vector, with $d$ denoting the number of action dimensions. Since actions directly control the system evolution, high-magnitude or poorly directed actions can result in constraint violations or unstable behaviors. To mitigate this, we introduce a regulator network $\rho_{\theta}:\mathcal{S}\times\mathcal{A}\times\mathbb{R}\rightarrow(0,1]^{d}$, which learns a scaling vector with an individual factor for each action dimension based on the current state, the raw action, and its predicted cost. At each step, the agent samples a raw action $a_{t}\sim\pi_{\phi}(\cdot|s_{t})$, computes the cost estimate: $\hat{c}_{t}=\max(Q_{c}^{1}(s_{t},a_{t}),Q_{c}^{2}(s_{t},a_{t}))$ from a twin-critic architecture, and the regulator network outputs a scaling vector $\rho_{t}$, see Fig. 2. The final action applied to the system is:
$\tilde{a}_{t}=\rho_{t}\odot a_{t},\quad\text{where }\rho_{t}=\rho_{\theta}(s_{t},a_{t},\hat{c}_{t}),$ (5)
where $\odot$ denotes element-wise multiplication, where each component of the action vector is multiplied by a scaling factor between 0 (high risk, large attenuation) and 1 (safe, no attenuation), smoothly reducing potentially unsafe actions proportional to predicted risk. This element-wise modulation attenuates each component of the action based on its risk profile, reducing the magnitude of high-risk components. Unlike hard safety constraints that may override agent behavior, this approach preserves the agent’s exploration behavior and allows stable off-policy learning. In many robotic domains, safety costs naturally increase with the magnitude of control inputs: high torques in manipulators accelerate wear and overheating, and large contact forces in legged robots risk joint damage. By designing costs that capture this structure, practitioners can align the regulator with system-specific safety considerations, making scaling an intuitive and broadly applicable mechanism for enforcing safety.
IV-C Learning Objectives and Updates
Reward Learning. We adopt a general off-policy reinforcement learning framework where the agent’s actor and critic are trained using the scaled action $\tilde{a}_{t}$, as this is the action that is actually executed in the environment. The reward critic is updated using:
$Q_{r}(s_{t},\tilde{a}_{t})\;\leftarrow\;$ $\;r(s_{t},\tilde{a}_{t})$
$+\gamma\,\mathbb{E}_{\begin{subarray}{c}s_{t+1}\sim p(\cdot|s_{t},\tilde{a}_{t})\\ a_{t+1}\sim\pi(\cdot|s_{t+1})\end{subarray}}\!\left[Q_{r}(s_{t+1},\tilde{a}_{t+1})\right],$ (6)
where $\tilde{a}_{t+1}=\rho_{t+1}\odot a_{t+1}$ and $\rho_{t+1}=\rho_{\theta}(s_{t+1},a_{t+1},Q_{c}(s_{t+1},a_{t+1}))$. The policy is updated to maximize the expected return under the regulated action:
$\mathcal{L}_{\text{actor}}=\mathbb{E}_{s_{t}\sim\mathcal{D},a_{t}\sim\pi(\cdot|s_{t})}\left[-Q_{r}(s_{t},\tilde{a}_{t})\right],$ (7)
ensuring that policy learning reflects the actual dynamics induced by the regulated action $\tilde{a}_{t}$. Our framework is algorithm-agnostic and can be integrated with any off-policy actor-critic method. For entropy-regularized algorithms such as SAC, the corresponding entropy term may be included in the actor objective. In our experiments, we demonstrate compatibility with both SAC and Twin Delayed DDPG (TD3)[28].
Cost Learning. The cost critic is also trained on the scaled actions using a TD-style Bellman backup [38]:
$Q_{c}(s_{t},\tilde{a}_{t})\;\leftarrow\;$ $\;c(s_{t},\tilde{a}_{t})$
$+\gamma\,\mathbb{E}_{\begin{subarray}{c}s_{t+1}\sim p(\cdot|s_{t},\tilde{a}_{t})\\ a_{t+1}\sim\pi(\cdot|s_{t+1})\end{subarray}}\!\left[Q_{c}(s_{t+1},\tilde{a}_{t+1})\right],$ (8)
This ensures the critic reflects the safety implications of the actual executed action $\tilde{a}_{t}$.
Regulator Objective. The regulator is trained to minimize the predicted cost of the executed action $\tilde{a}_{t}$, while avoiding degenerate solutions that collapse actions toward zero. Its loss function is given by:
$\mathcal{L}_{\text{reg}}=\;\mathbb{E}_{s_{t}\sim\mathcal{D},\,a_{t}\sim\pi(\cdot|s_{t})}\Big[$ $\beta\cdot Q_{c}(s_{t},\tilde{a}_{t})$ (9)
$-\lambda\cdot\log\rho_{\theta}(s_{t},a_{t},\hat{c}_{t})\,\Big].$
where $\beta,\lambda>0$ are trade-off parameters. The first term encourages the regulator to scale down actions that lead to high predicted costs. However, without the second term, a trivial solution where $\rho_{\theta}(s,a,\hat{c})\rightarrow 0$ would minimize this objective by collapsing all actions—halting the agent’s behavior entirely. To counteract this, the second term acts as a barrier penalty that diverges as any element of the scaling vector approaches zero. It encourages the regulator to retain as much of the original action magnitude as possible, unless high predicted cost necessitates suppression.
Algorithm 1 Cost-Aware Action Scaling (training loop)
1:Environment $\mathcal{E}$, replay buffer $\mathcal{D}$, actor $\pi_{\phi}$, regulator $\rho_{\theta}$, reward critic $Q_{r}$, cost critic $Q_{c}$
2:Initialize network parameters and target networks
3:for each interaction step do
4:  Observe $s$ and sample $a\sim\pi_{\phi}(\cdot\mid s)$
5:  $\hat{c}\leftarrow\max\!\big(Q_{c}^{1}(s,a),\,Q_{c}^{2}(s,a)\big)$ $\triangleright$ predicted cost
[…]
7:  Execute $\tilde{a}$, obtain $(r,c,s^{\prime})$, store $(s,\tilde{a},r,c,s^{\prime})$ in $\mathcal{D}$
8:end for
9:for each gradient step do
10:  Sample minibatch from $\mathcal{D}$
11:  Update $Q_{r}$ and $\pi_{\phi}$ via Eqs. (6)–(7)
12:  Update $Q_{c}$ (Eq. 8)
13:  Update $\rho_{\theta}$ (Eq. 9)
14:  Polyak-average target networks
15:end for
Optimality Trade-Off. The regulator’s training objective can be interpreted as solving a local constrained optimization problem at each state-action pair $(s_{t},a_{t})$:
$\min_{\rho\in(0,1]^{d}}\;\beta\cdot Q_{c}(s_{t},\rho\odot a_{t})-\lambda\cdot\log(\rho+\epsilon),$ (10)
where the logarithm is applied element-wise to the scaling vector $\rho$, and we include $\epsilon$ to avoid instability as $\rho\to 0$, ensuring gradients remain well-defined during training.
The coefficients $\beta$ and $\lambda$ balance the trade-off between minimizing predicted cost and preserving action magnitude: larger $\lambda$ encourages less suppression, while larger $\beta$ prioritizes cost reduction. Since $Q_{c}(s,\rho\odot a_{t})$ is typically a nonlinear function of the scaled action, the optimization problem lacks a closed-form solution but can be efficiently solved via gradient-based updates. This formulation ensures that the regulator selectively attenuates risky action dimensions while retaining as much of the agent’s original behavior as possible.
Gradient Flow and Modularity. To ensure clean modularity, we detach the scaling weights $\rho_{\theta}(s_{t},a_{t},\hat{c}_{t})$ from the computational graph when updating both the reward and cost critics, preventing gradients from flowing through the regulator. Similarly, the actor receives no gradients from the regulator, learning purely from task returns. Moreover, the regulator is updated independently via its own objective, ensuring that reward maximization and safety modulation remain decoupled. The full training procedure is summarized in Algorithm 1.
This design is particularly well-suited for off-policy reinforcement learning, where updates are performed using transitions stored in a replay buffer, independent of the current policy. Since the regulator modulates actions after sampling from the policy $\pi_{\phi}(\cdot\mid s)$, the executed action $\tilde{a}=\rho_{\theta}(s,a,\hat{c})\odot a$ differs from the originally sampled action $a$, and only the regulated action is stored and used for training. Off-policy methods naturally accommodate this, as policy and critic updates rely on the actual executed actions rather than the distribution used to generate them.
Implementation Details. The proposed method integrates a standard off-policy RL agent with a lightweight action regulator network for constraint satisfaction. The RL agent follows its baseline implementation without modification. The regulator is a feedforward neural network with two hidden layers of sizes [256, 256] and ReLU activations, outputting element-wise scaling factors between 0 and 1 via a sigmoid activation to modulate the agent’s actions. The regulator is trained using predicted costs from twin cost critics. Each cost critic is a feedforward network with two hidden layers of sizes [256, 256] and Tanh activations, taking state-action pairs as input and outputting a scalar cost prediction.
V Experiments
Figure 3: Performance comparison on the Safety Gymnasium locomotion environments. Each method is averaged over three independent runs; bold lines indicate the mean, and shaded areas show the standard deviation. Our methods (SAC-REG and TD3-REG) consistently achieve the best trade-off between return and cumulative constraint cost across all environments. Top: Episode return. Middle: Cumulative safety cost. Bottom: Return-to-Log-Cost ratio. Our methods outperform strong baselines, including PPO-PID [17], SIMMER [19], SRCPO [26], RESPO [25], and SEDITOR [24]. SIMMER is omitted from the Swimmer plot as it consistently yields negative returns, moving opposite to the target velocity.
Our experiments are designed to achieve the following objectives: (i) compare our approach against state-of-the-art safe RL baselines across different dynamical systems, (ii) analyze the influence of the key hyperparameters ($\lambda$ and $\beta$) from Eq. 10, which govern the trade-off between action preservation and cost suppression, (iii) evaluate the regulator’s action-scaling mechanism through ablations such as element-wise versus scalar regulation on different systems, and (iv) study robustness under injected sensor and actuator noise, highlighting the method’s potential for sim-to-real transfer.
Environments: We evaluate our method on locomotion tasks from the Safety Gym benchmark [21], namely Ant, Walker2d, Swimmer, HalfCheetah, and Humanoid. In these velocity tasks, a safety cost is incurred whenever the center-of-mass speed exceeds a predefined threshold. Because the cost signal is sparse and triggered only by such threshold violations, these environments provide a challenging setting for safe RL, while naturally aligning with our regulator’s goal of attenuating large actions that are most likely to induce violations. We further evaluate on BiGlucose and the Continuous Stirred Tank Reactor (CSTR) from the Safety-Critical Systems  [22].
Baselines: We compare our regulator against five state-of-the-art Safe RL baselines. PPO-PID [17] augments PPO with a PID-controlled Lagrangian multiplier to mitigate instabilities commonly observed in dual updates during constrained optimization. Simmer [19] augments PPO with a safety state that tracks the remaining safety budget. Safety Editor [24] uses two SAC agents, one for maximizing the reward and another for editing unsafe actions. RESPO [25] estimates reachability sets and constrains policy to remain within safe regions. SRCPO [26] formulates a bi-level constrained optimization using spectral risk measures to achieve a near-zero constraint violation rate while maximizing reward.
Metrics: Similar to prior safe RL studies, we report returns and cumulative costs as in [23, 25, 32], and additionally follow [23] in using the return-to-cost (RC) ratio to capture the trade-off between task performance and safety. Specifically, we measure (i) episodic return, (ii) cumulative cost during training, which reflects the total number of constraint violations and, in sparse-cost settings such as Safety Gym velocity tasks, implicitly captures violation frequency, and (iii) the RC ratio, defined as the total return divided by cumulative cost. For visualization, we plot the return divided by the logarithm of the accumulative cost, which improves interpretability of the safety-performance trade-off.
V-A Comparison Against Baselines:
Safety Gym Results
Across the locomotion tasks in the Safety Gymnasium suite, our methods—SAC-Regulator and TD3-Regulator—consistently deliver strong task performance while substantially reducing safety violations. Each method was evaluated over three random seeds; bold lines in Fig.3 denote the mean return across runs, with shaded regions representing standard deviation. Compared to the baselines, our approach achieves higher or comparable episode returns, indicating that soft action scaling does not hinder exploration. Notably, TD3-Regulator achieves the greatest cost reductions, with up to 126× lower cumulative cost in Walker2d, 64× in Ant, and 86× in Swimmer. Meanwhile, SAC-Regulator outperforms both TD3-Regulator and all baselines in HalfCheetah and Humanoid, achieving cost reductions of up to 28× and , respectively. RESPO can reach comparable returns when trained for 9M steps, but only with substantially higher violations and failure to converge in Humanoid.
Relative Return Improvement of SAC-Reg Over Baselines (↑)
Method Ant HalfCheetah Swimmer Walker2d Humanoid
PPO-PID [17] 27.33 3.04 0.32 3.76 7.13
SIMMER [19] 26.14 6.09 1.21 3.59 7.60
SEditor [24] 3.34 5.64 -0.10 1.95 18.52
RESPO [25] 23.97 2.89 0.17 204.40 12.73
SRCPO [26] -0.11 0.02 2.90 0.86 0.90
Relative Cost Compared to SAC-Reg (↓)
PPO-PID [17] 39.29 28.39 21.31 19.11 5.46
SIMMER [19] 41.88 13.70 54.22 14.14 3.64
SEditor [24] 16.19 1.67 1.77 0.48 1.88
RESPO [25] 8.36 1.65 27.39 2.03 1.29
SRCPO [26] 23.24 23.55 37.10 5.07 2.82
TABLE I: Relative return improvement (↑) and relative cumulative cost (↓) of SAC-Reg compared to baselines across locomotion tasks. SAC-Reg consistently achieves higher returns and lower cumulative costs than prior safe RL methods across all environments.
Relative Return Improvement of TD3-Reg Over Baselines (↑)
Method SafetyAnt HalfCheetah Swimmer Walker2d Humanoid
PPO-PID [17] 18.49 3.05 0.17 3.87 4.51
SIMMER [19] 17.70 6.11 1.18 3.70 4.83
SEditor [24] 2.55 5.65 -0.20 2.02 12.22
RESPO [25] 16.26 2.90 3.59 209.23 8.30
SRCPO [26] -0.41 0.02 2.46 0.91 0.29
Relative Cost Compared to TD3-Reg (↓)
PPO-PID [17] 60.14 19.77 34.06 126.18 5.20
SIMMER [19] 64.11 9.54 86.65 93.37 3.47
SEditor [24] 24.78 1.16 2.84 3.15 1.74
RESPO [25] 12.80 1.15 43.78 13.42 1.23
SRCPO [26] 35.57 16.39 59.28 33.50 2.69
TABLE II: Relative return improvement (↑) and relative cumulative cost (↓) of TD3-Reg compared to baselines across locomotion tasks.TD3-Reg demonstrates similar trends, outperforming baselines in return while maintaining substantially lower cumulative costs.
Tables I and II summarize results for both regulators against established baselines. Two metrics are reported: the relative return improvement,
$\frac{\text{Return}_{\text{Ours}}-\text{Return}_{\text{Baseline}}}{|\text{Return}_{\text{Baseline}}|},$
and the relative cumulative cost of each baseline normalized by our method. Positive return values indicate improved task performance, while cumulative cost ratios above 1.0 indicate higher constraint violations than ours. For example, in Walker2d, SAC-Reg outperforms RESPO with a relative return improvement of 204.4, corresponding to a 20,440% increase.
Overall, our methods deliver the lowest cumulative cost across all tasks without compromising return. Unlike approaches such as SEditor or RESPO, which improve safety at the expense of performance, our regulators preserve exploration and consistently achieve superior return-to-cost trade-offs. This demonstrates the effectiveness and generality of decoupling safety regulation from reward learning.
Experiments on Safety Critical Systems
Figure 4: Performance comparison on the BiGlucose and CSTR environments. Our method (SAC-Reg) achieves high return with low cost, yielding the best return-to-cost ratio throughout training.
Figure 5: Ablation Study for evaluating the impact of the regulator hyperparameters $\lambda$ and $\beta$ on Return, Cumulative Cost, and Return-to-Cost ratio. Each curve shows the mean across three runs, and shaded regions indicate standard deviation. The top row varies $\lambda$ with fixed $\beta=10$; the bottom row varies $\beta$ with fixed $\lambda=0.0015$. As seen, smaller $\lambda$ values reduce cumulative cost, with $\lambda=0.0015$ giving the best balance between performance and safety, while $\beta=10$ provides the most favorable trade-off overall.
Figure 6: Comparison between element-wise and scalar action scaling. While both perform similarly in most tasks, the scalar variant fails to converge in the high-dimensional Humanoid environment, indicating that element-wise scaling improves safety and stability in complex control settings.
We evaluate our approach on two continuous-control environments: BiGlucose and the Continuous Stirred Tank Reactor (CSTR) from [22], which capture biomedical and chemical process dynamics. BiGlucose models blood glucose regulation with insulin and glucagon injections under delayed, partially observable dynamics, requiring glucose levels to stay within physiological bounds. CSTR simulates nonlinear reactor dynamics where unsafe control can cause hazardous runaway reactions. In both cases, we modify the environments to provide a continuous cost signal by logging violation magnitudes, instead of terminating episodes on constraint breaches. For full specifications, see [22].
Figure 4 compares our SAC-Reg method against baselines (SEditor, SIMMER, SRCPO, and PPO-PID). All methods are trained for the same number of environment steps as in [22]; RESPO is excluded since it takes significantly more steps to converge. Across both tasks, SAC-Reg consistently achieves higher returns with fewer cumulative constraint violations, yielding superior safety–performance trade-offs. In BiGlucose, baselines such as SIMMER and SRCPO quickly accumulate costs despite improving return, while SEditor remains return-limited. In CSTR, only our method manages both safety and performance, as others either accumulate high violations or fail to learn.
Ablation Study on $\lambda$ and $\beta$
We conduct an ablation study in the Ant environment to evaluate the sensitivity of our regulator framework to the hyperparameters $\lambda$ and $\beta$, which control the trade-off between action retention and cost suppression. When varying $\lambda$ over the range $\{1\times 10^{-5},\ 0.0015,\ 0.05,\ 0.25,\ 1.0\}$, we find that smaller values lead to significantly lower cumulative costs. In particular, $\lambda=0.0015$ achieves the best balance between constraint satisfaction and task performance. Higher values of $\lambda$ result in larger action magnitudes and consequently higher constraint violations. Similarly, varying $\beta$ over $\{5,\ 10,\ 15,\ 30,\ 50\}$ shows that $\beta=10$ achieves the best overall safety-performance balance, minimizing constraint violations while maintaining high return. These results, as shown in Fig. 5, highlight the importance of properly tuning the regulator’s loss coefficients to achieve optimal return-to-cost behavior.
Element-wise vs. Scalar Regulation
To evaluate the impact of element-wise action regulation, we conducted an ablation study comparing our full regulator with a simplified variant that uses a single scalar value to uniformly scale all action dimensions. Figure 6 presents results across all Safety Gymnasium locomotion tasks. While the scalar variant achieves comparable performance in most environments, it fails to converge in the high-dimensional Humanoid task. This suggests that element-wise scaling is particularly important in complex control settings, where individual action dimensions exhibit distinct risk profiles. Fine-grained modulation allows the regulator to target risky joints more precisely, improving both safety and learning stability.
Step Noise 0.00 Noise 0.025 Noise 0.05 Noise 0.10
Return Cost Return Cost Return Cost Return Cost
100k 1961 6 753 0 762 28 677 1
300k 2558 79 1554 78 1614 165 1103 113
500k 2492 199 1945 195 1969 282 1381 387
700k 2533 282 2139 230 2071 339 1596 572
900k 2501 388 2104 262 2074 479 1589 767
1000k 2571 412 2016 312 2089 533 1510 849
TABLE III: Training performance under different levels of injected Gaussian noise ($\sigma=0.00,0.025,0.05,0.10$) in observations and actions. Values show episode return and cumulative cost at checkpoints. Our regulator maintains bounded costs across noise levels, demonstrating robustness relevant for sim-to-real transfer.
Robustness and Sim-to-Real Transfer
To approximate uncertainties encountered on physical robots, we inject Gaussian noise into both observations and actions during training, modeling sensor measurement errors and actuator execution noise. Agents are trained with noise levels ($\sigma=0,0.025,0.05,0.10$), and the resulting training performance is summarized in Table III. Across all noise settings, our regulator achieves strong returns while keeping cumulative costs bounded. Even under the highest noise level ($\sigma=0.10$), performance remains stable, highlighting robustness to sensing and actuation imperfections and supporting the method’s potential for sim-to-real transfer.
VI Conclusion
We introduced a modular and practical framework for safe reinforcement learning that decouples reward maximization from safety enforcement through a cost-aware regulator. Instead of overriding agent actions, our method scales them smoothly based on predicted constraint violations, preserving exploration and enabling stable off-policy learning. The regulator uses twin cost critics for robust cost estimation and is trained with a loss that balances risk reduction and action preservation. Our approach is model-free and integrates seamlessly with existing off-policy RL pipelines. Empirical results on diverse benchmarks demonstrate that our method consistently achieves the highest return-to-cost ratios, reducing constraint violations by up to 126 times while maintaining or improving task performance relative to prior state-of-the-art methods. The regulator aligns with real-world safety limits such as torque bounds in manipulators, and joint load management in legged robots. Robustness experiments with injected observation and action noise further demonstrate bounded costs and stable returns under uncertainty, supporting the potential for sim-to-real transfer. A key direction for future work is to develop principled strategies for automatically tuning the regulator hyperparameters ($\lambda$ and $\beta$) and to extend the approach beyond input-magnitude costs toward more general safety constraints.